# AI sustains higher strategic tension than humans in chess

### v1

@@ I Defining Tension in Chess
While engines can maintain higher tension for a longer duration compared to humans, how does this difference depend on skill level? We consider the cumulative tension up to ply 150 across different Elo ratings for human players in Figure 5a  [^3] In principle, Stockfish has a search depth parameter that serves as a proxy for skill level. Leela has a time limit parameter, during which it is permitted to assess the state of the board and make a decision. At a few dozen search depths, the tension load is limited to about 450, similarly for Leela at a few seconds. Since the games that are shown go far beyond these limited parameter values with supercomputers, this suggests that algorithmic complexity increases the ability to handle tension.. Tension changes substantially with Elo, revealing discrete jumps that separate human players into roughly three, significantly separate groups. A plateau in tension load from Elo of 1000 ends between 1500 and 1700. This transition typically corresponds to the entry point into competitive chess, where players begin to study openings, apply strategic principles more consistently, and reduce the frequency of major tactical blunders. After a second plateau, a jump in tension occurs at 2100 Elo. This point marks the entry into the realm of stronger players, such as experts and grandmasters, whose games exhibit more nuanced, complex positions with deeper strategic and tactical layers.
Figure 5b shows the tension load values for various fixed search depths of Stockfish. These results indicate an approximately constant tension load across different depths, suggesting a qualitatively different scenario compared to the Stockfish vs. Leela games from the TCEC tournaments (highlighted as the red point in Figure 5a), where the engines are free to adjust their search depth and, more generally, the time allocated per move  [^4] Representing the dependence of tension on the time limit per move in the case of Leela is challenging because it strongly depends on the technical specifications of the machine used, particularly the presence of a suitable dedicated GPU and sufficient RAM. In particular, achieving Elo ratings significantly higher than the best human players requires substantial hardware resources.. The correlation between Stockfish search depth and Elo rating is not well established. Previous experiments on a different engine (Houdini) reported a roughly linear relationship up to a search depth of 20 plies [13].
II Discussion
We investigate the dynamics of strategic tension, a network-based definition of piece interactions, in chess games played by humans and AI. Our central hypothesis is that distinct tension profiles would emerge, reflecting differing strategic approaches. The results confirm this hypothesis and reveal several key insights. First, we observed that state-of-the-art chess engines (Stockfish at high search depths and Leela Chess Zero at advanced neural network settings) consistently maintain significantly higher levels of strategic tension for longer durations than even the best grandmasters. This suggests an ability to navigate and sustain complex, balanced positions where multiple pieces exert pressure and threats remain unresolved. Second, human players exhibit a notable relationship between skill level (Elo rating) and tension management. Significant differentiation in accumulated tension emerges primarily above an Elo rating of approximately 1600, coinciding with the stage where players typically develop deeper strategic understanding, theoretical knowledge, and enhanced calculation abilities [9, 7]. This implies that managing higher levels of strategic tension is characteristic of advancing chess expertise in humans. Likewise, one might expect greater computational complexity to allow an engine to find and sustain even more complex positions, leading to a monotonic increase in tension load with search depth, similar to the trend observed with human expertise. Preliminary calculations indicate that the computational complexity of Stockfish, as reflected in its search depth, does influence tension load, but in a nonlinear manner: tension load exhibits flat plateaus across a wide range of fixed search depths (approximately from 10 to 30). This raises the question of how and where new play styles emerge as the complexity accessible to the algorithm increases.
The observed differences in tension dynamics distinguish how biological and artificial systems approach problem-solving under strategic constraints. Human cognition appears adapted to manage, and perhaps even actively reduce, high levels of sustained complexity, possibly due to inherent limitations in working memory, attention, or susceptibility to psychological pressure [15, 18]. State-of-the-art AI, optimized for performance within defined rule sets, demonstrates a capacity for tolerating and exploiting complexity [30]. The superior ability of high-level engines to sustain tension elucidates their ability to handle immense combinatorial complexity [34, 22]: engines do not simply calculate better but actively maintain more complex strategic states. This aligns with the idea that optimal strategy might be contingent on the agent’s cognitive or computational architecture. More broadly, such differences may be essential for AI systems designed for human interaction. While direct extrapolation requires caution, we might think of how to examine tension in other complex adaptive systems, such as economic markets or geopolitical negotiations [32]. The finding that increased computational complexity allows AI to sustain higher tension may matter for stability in international relations, potentially enabling longer standoffs but also increasing the stakes should tension break [14, 31] in an AI-enabled analogy to Dr. Strangelove.
@@ II Discussion
Several avenues for future research emerge from this work. The network-based tension metric represents one specific operationalization of strategic complexity. It primarily captures piece interactions and may not fully encompass other crucial factors like pawn structure, king safety, or long-term positional plans. While it changes predictably with several measures of board complexity (Result 2), parallel measures may help inform pedagogical approaches in chess. Further investigations could also focus on how different engines manage tension, examining variations across engine architectures, versions, and underlying algorithms. Extending this methodology to other complex strategic games, such as Go [36], Shogi, or even real-time strategy e-sports, could test the universality of the observed dynamics across different game structures (e.g., imperfect information, different spatial dynamics). Directly investigating the cognitive and physiological correlates of our tension metric in human players—for example, by combining game analysis with eye-tracking, think-aloud protocols, or stress indicators (like galvanic skin response)—could help elucidate the underlying mechanisms driving human tension. Furthermore, exploring whether AI can be explicitly trained to modulate a tension profile, perhaps to mimic human playstyles or achieve specific pedagogical goals, represents an interesting direction for human-AI interaction research. Investigating the relationship between our tension metric and game outcomes (win/loss/draw) beyond simple averages could also reveal critical thresholds or patterns predictive of game resolution. Finally, our analysis relies on existing game databases. While large, these datasets may have inherent biases, and game conditions (e.g., time controls, stakes) might differ from controlled experimental settings. The specific AI engine used (Stockfish) represents one architecture; other engines with different algorithms (e.g., neural network-based like AlphaZero beyond its initial papers) might exhibit different tension dynamics. While this study focused on Stockfish and Leela due to their prominence and availability, the analytical framework and open-source code we provide can be readily applied to other UCI-compatible engines, paving the way for broader comparative studies.
We find evidence for distinct approaches to managing strategic tension between human players and AI chess engines. By proposing a network-based metric, we characterize how tension typically accumulates and dissipates throughout games, revealing that high-performance AI sustains significantly greater complexity for longer periods than even elite human players. Human tension management appears correlated with expertise but potentially constrained by cognitive factors, while AI’s capacity in tension management is primarily a function of computational resources, which allows it to consider high-dimensional configuration of balanced interactions. While human players would find such algorithmic complexity inaccessible, the brain is the outcome of a trade-off in terms of longer-term evolutionary games. The short-term cost of losing a chess game might be worth the long-term benefits of metabolic cost and cognitive plasticity in the game of life.
@@
AC developed the code, performed the numerical analyses, and generated the figures. VDPS provided the original inspiration for the project and supervised its development at all stages. EDL participated in discussion. All authors contributed to writing the manuscript.
@@ III.1 Data
These datasets served as the basis for the analyses presented in Figure 2.
In addition, a large set of games played by human players with Elo ratings ranging from 1100 to 2300 was downloaded from https://www.ficsgames.org/, an online archive that collects games played on the Free Internet Chess Server (FICS). This dataset broadly represents amateur to intermediate-level players, allowing for analysis across a wide skill spectrum.
Together with the previously described datasets, these data formed the basis for the analyses presented in Figure 5.
III.2 Software
@@
III.3 Generating Stockfish games at varying depths
@@
IV Supplementary materials
IV.1 Why the spectral radius as a measure of tension
Topological entropy, $h_{\mathrm{top}}$, is a concept from dynamical systems theory that quantifies the exponential growth rate of the number of distinguishable orbits or, in the context of graphs, the diversity of paths. For unweighted networks (where the adjacency matrix has entries in $\{0,1\}$), the spectral radius $\lambda_{1}$ of the adjacency matrix is directly related to topological entropy. In particular, it has been shown that $\lambda_{1}$ satisfies a variational principle equivalent to the dynamical entropy of the network (i.e., the entropy of a stochastic process on the graph), leading to the definition of topological entropy as
$h_{\mathrm{top}}=\log(\lambda_{1}).$
Topological entropy was first introduced in 1965 by Adler, Konheim, and McAndrew [1] as a measure of the complexity of a dynamical system. In the context of networks, this relationship is particularly meaningful: the quantity $\lambda_{1}$ does not merely reflect the number of walks in the graph, but its logarithm measures the average Shannon information per step gained when observing a random walk unfold. A larger $\lambda_{1}$ indicates a greater number of possible distinct paths at each step, implying higher unpredictability, richer structural diversity, and overall complexity.
This connection positions $\lambda_{1}$ as a fundamental quantity bridging structural graph theory, information theory, and dynamics. It can be interpreted as a proxy for the “intrinsic computational capacity” or “information-processing potential” of the network.
In our study, this interpretation is particularly useful: we use the spectral radius $\lambda_{1}$ as a global measure of the tension and complexity present on the chessboard. A higher $\lambda_{1}$ suggests a denser and more interconnected pattern of potential interactions between pieces (e.g., threats and defenses), and therefore a richer position in tactical possibilities and strategic depth. This justifies using $\lambda_{1}$ as a meaningful descriptor of the game’s complexity at a given state.
IV.2 Maximized tension game
If we consider the unrealistic and absurd game in which, at each ply, the players randomly choose one of the moves that maximizes tension, by ply 52 the tension reaches a value of $T\simeq 10.3$ (Figure 6).
Figure 6: Tension network obtained by maximizing tension move by move. (a) Chessboard at ply 44, obtained by choosing the legal move that increases the tension the most at each ply. (b) The corresponding network represents the interactions on the chessboard, using the same color and edge scheme as in Figure 1.
The first 60 moves of the game are listed below: 1. d3 h6 2. Bg5 e5 3. Be7 Nc6 4. Qd2 d5 5. Qxh6 Qd6 6. Qe6 Nd8 7. c4 Rh6 8. cxd5 e4 9. Nc3 g5 10. Na4 b5 11. Rc1 Rb8 12. Rxc7 Rb7 13. Nc5 Qxd5 14. dxe4 Rb6 15. e3 Ra6 16. Rxa7 b4 17. Bd3 Bg7 18. Nf3 Bf8 19. Nxg5 Bd7 20. f4 Bc8 21. Kf2 b3 22. Rc1 bxa2 23. Rc3 a1=Q 24. Ra3 Qd1 25. Ra4 Qg4 26. Be2 Qgf5 27. Bd3 Qfxe4 28. Ra8 Nxe7 29. Kg3 Rh4 30. Kf2 Rh6.
After ply 52, the game maintains a relatively constant level of tension indefinitely, indicating that the game ceases to progress meaningfully.
IV.3 Stockfish evaluation for humans and engines
To better analyze the progression of games, we also examined the Stockfish evaluation, again distinguishing between games played by humans or AIs, and whether they ended in a draw or not (Figure 7). What emerges is that games between AIs show a rapid increase in the absolute value of the Stockfish evaluation from the very first plies. As the game progresses, the average evaluation values begin to diverge depending on the eventual outcome: games that end in a win or loss exhibit a steady increase in evaluation with each ply, whereas those that result in a draw maintain relatively low evaluation values throughout.
In contrast, in human games, the evaluation remains near zero during the opening phase. It is worth noting that, in fact, only starting from 2023 did official rules begin allowing Leela Chess Zero to select its openings directly from a database. However, as the game progresses, a similar divergence in evaluation can be observed. Notably, this shift typically occurs around ply 40, where tension generally peaks, marking a transition where evaluations for decisive games begin to rise consistently. In contrast, those for drawn games remain comparatively stable. Barthelemy [2] has already observed this result, using a similar framework for capturing the notion of fragility within chess games.
In the case of human games, the average values of tension are characterized by greater fluctuations, likely due to the lower predictability of human play and, more specifically, the greater variability between individual games, especially when different players are involved. Moreover, when focusing on tension values associated with drawn games, a notable difference emerges: in the case of AI matches, tension decreases towards zero much more gradually. This is mainly because engine games are longer on average than human games.
Figure 7: Stockfish evaluation during games for humans and AI. We consider 120 games for each (a subset of those analyzed in Figure 2), with evaluations computed at a fixed depth of 20. The same color scheme described in Figure 2 is used: cold and warm colors for humans and AI, respectively, and light/dark shades to distinguish draws from decisive games. We also show the overall average evaluations, along with vertical dashed lines indicating the ply corresponding to the maximum tension for both groups, humans and AI, shown in blue and red, respectively.
To better understand the type of information conveyed by the Stockfish evaluation, we analyzed the 2,400 games to compute the probability that a player goes on to win the game when holding an evaluation advantage of at least 100 centipawns at ply 40. In the case of human games, this probability is approximately 71 %, whereas for AI games, it is around 59 %.
IV.4 Additional properties of the chessboard and network
In this section, we present further comparative metrics between human and AI chess games, including the number of pieces, the sum of the weighted material in the chess board - where pieces are valued at 1 (pawns), 3 (knights and bishops), 5 (rooks), 9 (queens), and 0 (kings) - along with the number of links and the number of loops in the tension network. These measures provide complementary insights into positional advantages and dynamic board conflicts across different types of games (Figure 8).
IV.5 Correlation between tension and the number of chess pieces
We investigated the correlation between the measured tension values and the number of chess pieces present on the board. Naturally, this correlation is highly dependent on the stage of the game being considered. For this reason, we analyzed three distinct moments in the game—specifically at ply 30, 60, and 90—to show how, on average, tension varies with the number of pieces on the board, both for the AI and for human players (Figure 9).
We observe that, on average, tension increases with the number of chess pieces across all three stages of the game. However, this increase diminishes rapidly as the number of pieces grows, eventually reaching an approximately constant behavior—or even a slight decrease. This may be because when the number of pieces is significantly higher than what would typically be expected at a given ply, it might indicate that no major confrontation has yet occurred. Such confrontations often lead to a rise in tension, as more pieces are brought into key positions on the board. Moreover, an excessive number of pieces on the board can severely limit their mobility, reducing their ability to control important areas or pose threats to the opponent. Therefore, it is reasonable to expect that, to maximize tension, there exists an ideal range for the number of pieces — neither too high nor too low.
Figure 8: Comparison of additional structural properties across human and AI games. We analyze the set of games shown in Figure 2. Four measures are considered: the number of pieces, the weighted material, the number of links, and the number of loops in the interaction network. The color scheme is red and blue for AI and human games, respectively.
Figure 9: Analysis of the dependence of tension on the number of chess pieces at three game moments. We analyze the set of games shown in Figure 2. Average tension values are shown for each number of chess pieces, distinguishing between three game moments: ply 30, 60, and 90. The color scheme is red and blue for AI and human games, respectively. The error bars correspond to the standard deviation calculated for each represented value.

### v2

@@ II Result 2: Strategic tension reflects, does not predict game result
While engines can maintain higher tension for a longer duration compared to humans, does this difference depend on skill level? We consider the cumulative tension up to ply 150 across different Elo ratings for human players in Fig. 5a  [^3] In principle, Stockfish has a search depth parameter that serves as a proxy for skill level. Leela has a time limit parameter during which it is permitted to assess the board state and make a decision. At a few dozen search depths, the tension load is limited to about 450, similarly for Leela at a few seconds. Since the games that are shown go far beyond these limited parameter values with supercomputers, this suggests that algorithmic complexity increases the ability to handle tension.. Tension changes substantially with Elo, showing an approximately linear dependence in human games. Moreover, across all Elo values, rapid games consistently exhibit slightly higher tension loads than blitz games. This is consistent with the idea that longer time controls allow human players to create and manage more complex positions [5]. The effect is even more pronounced in classical games, where players have significantly more time to make decisions, resulting in higher tension and reflecting the chess community’s concern that time-constrained players experience larger performance drops on complex decisions and therefore tend to choose safer, less strategically risky continuations [5, 7]. In contrast, elite classical play is often characterized by keeping simplification at bay and maintaining a complex struggle, where grandmasters sustain tense, fluid positions that test the opponent’s stamina and depth of calculation over many hours [22, 30, 12].
For comparison, we examine how the tension load changes for Stockfish, where we can control the computational complexity of the algorithm by adjusting the search depth at a fixed value for simulated games in Fig. 5b. Although we examined search depths only up to 50, the results already indicate that increasing depth leads to higher tension. As a reference, previous experiments using a different engine (Houdini) reported a roughly linear relationship between Elo rating and search depth up to 20 plies [17]. Overall, this indicates that increased tension load is indicative of greater computational demands on the player.
III Discussion
We investigate the dynamics of strategic tension, a network-based definition of piece interactions, in chess games played by humans and AI. Our central hypothesis is that distinct tension profiles would emerge, reflecting differing strategic approaches. Our observations confirm this hypothesis and reveal that state-of-the-art chess engines (Stockfish at high search depths and Leela Chess Zero at advanced neural network settings) consistently maintain significantly higher levels of strategic tension for longer durations than even the best human grandmasters.
We speculate that the sustained tension in AI play reflects the deferral of immediate opportunities and facilitates latent strategic options that only materialize in later phases of the game. This kind of strategic decision making would be facilitated by the ability to handle greater computational complexity. Correspondingly, we find that human players exhibit a linear relationship between cumulative tension and skill level (Elo rating). Similarly, Stockfish’s tension load increases with computational complexity, as reflected in its search depth. Far beyond these examples are AI engines running on supercomputers. Superhuman computational capacity obviates the need to simplify upon leaving theoretical territory; AI engines do not need to rely on heuristics and intuition such as familiarity with positions in opening theory that heavily guide expert human play. These observations suggest that the ability to sustain strategic tension is a hallmark of (for human players) advanced chess expertise, encompassing deeper strategic understanding, and extensive theoretical knowledge and (for AI) enhanced calculation skills [12, 9].
The increase in tension with computational load likely arises from different factors in the two groups, reflecting their nature as artificial versus biological systems. In AI, tension rises as engines gain computational capacity and play at higher levels, where controlling more pieces and squares maximizes options and exploits opponents’ mistakes [4]. In humans, more skilled players can handle complex positions and maintain higher‑tension situations without simplifying [23, 11, 10, 18]. Indeed, humans may tend to simplify under certain conditions possibly due to limitations in working memory, attention, or susceptibility to psychological pressure [20, 26]. This interpretation of the results could also suggest that human players may sometimes actively steer games toward drawn resolutions, reflecting deliberate risk minimization and metagame considerations typically absent in AI play. The simplification of high-tension positions thus serves not only strategic purposes but also as an adaptive response to the computational demands of accurate tactical calculation, allowing players to sustain the level of accuracy needed for strong play within cognitive limits [14]. Of course, such limits come from the evolutionary trade-offs reflected in the brain, where short-term costs, such as losing a chess game, are outweighed by long-term benefits in metabolic efficiency and general, rather than domain-specific, intelligence [snell-roodBrainSizeGlobal2009, burgerAllometryBrainSize2019].
@@ III Discussion
Several avenues for future research emerge from this work. Extending this methodology to other complex strategic games, such as Go [38], Shogi, or even real-time strategy e-sports, could test the universality of the observed dynamics across different game structures (e.g., imperfect information and varying spatial dynamics). Moreover, the AI engine used (Stockfish) represents one architecture; other engines with different algorithms, such as neural network-based engines like AlphaZero, may show different tension dynamics. While we focused on Stockfish and Leela due to their prominence and availability, our analytical framework and open-source code can readily be applied to other UCI-compatible engines, enabling broader comparative studies. With a more direct focus on tension, directly investigating the cognitive and physiological correlates of our tension metric in human players—for example, by combining game analysis with eye-tracking, think-aloud protocols, or stress indicators such as galvanic skin response—could help elucidate the underlying mechanisms driving changes in tension.
The observed differences between artificial and human systems could have implications for AI designed to interact with humans, particularly as AI is expected to play an increasing role in teaching and training across many domains. In chess, a notable step in this direction is the Maia engine [29], which is explicitly designed to model human play rather than engine-optimal moves. Moreover, while direct extrapolation requires caution, examining how tension is sustained in other complex adaptive systems, such as economic markets or geopolitical negotiations, can provide insight [35]. The finding that increased computational complexity leads AI to sustain tension may matter for stability in international relations [simmons-edlerPositionAIPowered2024], potentially enabling more complex standoffs but also increasing the stakes of miscalculation [19, guttieriIntegrativeComplexity1995, 34], creating an AI-enabled parallel to Dr. Strangelove.
@@
AC led the definition of the tension metric, with contributions from the other authors, and carried out the code development, numerical analyses, and figure generation. EDL performed several simulations on a computing cluster using Stockfish at high depths. VDPS provided the original inspiration for the project, supervised its development at all stages, and independently cross-checked the results. All authors participated in the discussions and contributed to writing the manuscript.
@@ IV.1 Data
These datasets served as the basis for the analyses presented in Fig. 2.
In addition, a large set of rapid and blitz games played by human players with Elo ratings ranging from 1000 to 2700 was downloaded from https://lichess.org/, an online platform that provides a vast archive of games. This dataset broadly represents amateur to expert players and is ideal for analysis because games are subject to anti-cheating controls, ensuring reliable data across a broad skill spectrum.
Together with the previously described datasets, these data formed the basis for the analyses presented in Fig. 5.
IV.2 Software
@@
IV.3 Generating Stockfish games at varying depths